A STUDY OF NbMo AND NbMo^{−} BY ANION PHOTOELECTRON SPECTROSCOPY

\underline{PRAVEENKUMAR BOOPALACHANDRAN}, SRIJAY S. RAJAN, MELISSA A. BAUDHUIN, and DOREEN G. LEOPOLD , Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

We report the 488 and 514 nm anion photoelectron spectra of NbMo^{−}. R2PI spectroscopic studies have established that neutral NbMo has a ^{2}Δ_{5/2} ground state and a short bond length ^{a}. We find that the NbMo^{−} anion has a ^{1}Σ^{+} ground state in which the ”extra” electron occupies the (4d)δ bonding orbital, giving a 1σ^{2}1π^{4}1δ^{4}2σ^{2} valence electron configuration. Thus, NbMo^{−} has a formal bond order of 6, and is isoelectronic with Mo_{2}. Low-lying excited states of NbMo (^{2}Σ^{+}) and NbMo^{−} (^{3}Δ) are also observed. The spectra provide the electron affinity of NbMo, energies of the ^{2}Σ^{+} and ^{3}Δ excited states, vibrational frequencies in the anion and neutral molecule ground states and the ^{2}Σ^{+} state, ^{2}Δ and ^{3}Δ spin-orbit splittings, and (from Franck-Condon analyses) bond length measurements for the anion ground state and the observed excited states. These results are compared with previous anion photoelectron spectroscopic data for the Group 5/6 congeners VCr and VMo, and with density functional theory predictions.

aR. Nagarajan and M. D. Morse, J. Chem. Phys. 127, 164305 (2007).